OpenStudy (anonymous):

what are the factors of x^2-5x+6 ?

jimthompson5910 (jim_thompson5910):

What two numbers multiply to 6?

OpenStudy (anonymous):

3 & 2

jimthompson5910 (jim_thompson5910):

notice how they add to 3+2 = 5 we want this sum to be -5 and NOT +5

OpenStudy (anonymous):

correct

jimthompson5910 (jim_thompson5910):

however, -3 + (-2) = -5 and -3 * (-2) = 6 so the two numbers -3, -2 multiply to 6 and add to -5

jimthompson5910 (jim_thompson5910):

meaning that x^2-5x+6 factors to (x-3)(x-2)

OpenStudy (anonymous):

so how would i solve the problem

jimthompson5910 (jim_thompson5910):

it's shown above

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} x^2&-5x&+6\\ &\uparrow &\uparrow \\ &-3-2&-3\cdot -2 \end{array}\)

OpenStudy (anonymous):

jim_thompson5910 how would i do it step by step solving

jimthompson5910 (jim_thompson5910):

it's more like a trial and error kind of thing you list out the possible ways to multiply to 6. Then you add up the factors to see which pair adds to -5

jimthompson5910 (jim_thompson5910):

6 = 6*1 6+1 = 7... doesn't work 6 = 3*2 3+2 = 5 ... doesn't work etc etc

OpenStudy (anonymous):

would you do the solving i am stuck

jimthompson5910 (jim_thompson5910):

it's already solved for you

OpenStudy (anonymous):

where

OpenStudy (aum):

You were asked to factor x^2-5x+6 and jim-thompson showed you it factors to: (x-3)(x-2). Those are the two factors.

OpenStudy (aum):

x^2 - 5x + 6 = x^2 - 2x - 3x + 6 = x(x-2) - 3(x - 2) = (x-2)(x-3)

OpenStudy (anonymous):

That the process ?

OpenStudy (aum):

Yes. But look at jim_thompson's explanation on how to arrive at the -2 and -3 numbers. In x^2-5x+6 you will have to ask yourself what two numbers multiply to +6 and add to -5? Here, -2 and -3 multiply to +6 and add to -5. Once we know the two numbers we can straightaway write the factors as (x-2)(x-3) or if you prefer you can proceed in steps as shown in my previous reply.

OpenStudy (aum):

If you wish you can try factoring x^2 + 3x - 10

OpenStudy (anonymous):

Correct but how can i first solve im really stuck

OpenStudy (aum):

Are you asked to factor or are you given an equation to solve?

OpenStudy (anonymous):

factor

OpenStudy (aum):

Let us try factoring x^2 + 3x - 10. The last term is -10 and the middle coefficient is +3 What two numbers multiply to -10 and add to 3?

OpenStudy (anonymous):

5 and 2

OpenStudy (aum):

Correct. But +5 and -2 because 5 * (-2) = -10 (last term) 5 + (-2) = 3 (middle coefficient) So the factors are (x+5)(x-2)

OpenStudy (anonymous):

Oh okay so to factor x^2- 5x =6 is what ?

OpenStudy (anonymous):

3 and 2 correct

OpenStudy (aum):

First bring the 6 to the left: x^2 - 5x - 6 = 0 You might be tempted to say 3 and 2 but to multiply 3 and 2 and get a -6 they have to be of opposite signs. If you try various combinations: -3 and + 2 or 3 and -2 it does not add to -5. So consider some other factors of 6 other than 3 and 2. What two other numbers multiply to 6?

OpenStudy (aum):

How about 1 and 6? They multiply to two and with proper sign they can be made to add to -5. 1 and -6 multiply to -6 1 + (-6) add to - 5 So the factors are: (x+1)(x-6)

OpenStudy (anonymous):

Oh

OpenStudy (aum):

Yes. Try various factors. If one set of factors does not work try another set.

OpenStudy (aum):

How about x^2 - 5x - 36?

OpenStudy (aum):

Factors of 36 are: 1 and 36 2 and 18 3 and 12 4 and 9 6 and 6 Which one of these pairs, with proper sign would multiply to -36 and add to -5?

OpenStudy (anonymous):

12 and 3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

please factor my problem

OpenStudy (aum):

which one? All the problems you posted have been factored.

OpenStudy (anonymous):

Im just asking for x^2-5x+6

OpenStudy (aum):

But earlier you posted it as x^2- 5x =6 ( '=' instead of '+'). Typo?

OpenStudy (anonymous):

sorry it was a typo

OpenStudy (anonymous):

its a plus sign

OpenStudy (aum):

x^2-5x+6 Same question: What two numbers multiply to +6 and add to -5?

OpenStudy (anonymous):

3 and 2

OpenStudy (aum):

Yes, 3 and 2 is correct. Next, what should be the signs of 3 and 2. Remember they have to multiply to +6 which means both have to be the same sign. Either +3 and +2 OR -3 and -2. In both cases they will multiply to +6. But it should also add to -5. So what should be the signs? +3 and +2 OR -3 and -2?

OpenStudy (anonymous):

3 is negative and 2 positive

OpenStudy (aum):

They cannot be opposite signs because they will multiply to a negative number but we want positive 6. So they have to be the same sign. Both positive or both negative.